# T2VEval: Benchmark Dataset and Objective Evaluation Method for T2V-generated Videos

### v6

@@ 1 Introduction
Text-to-video (T2V) technology has made significant advancements in the last two years and attracted growing attention from both the research community and the general public. As an important branch of Generative Artificial Intelligence (GenAI), T2V models Blattmann et al. (2023b, a); Ho et al. (2022); Singer et al. (2022); Wu et al. (2021, 2022a); Hong et al. (2022); Villegas et al. (2022); Zhou et al. (2022); Khachatryan et al. (2023); Luo et al. (2023); He et al. (2022); Wang et al. (2023); Yang et al. (2024); Zhang et al. (2024a); Wang et al. (2024); Brooks et al. (2024) have developed rapidly recently. Contemporary T2V models, such as Sora, Runway Gen-3, PixVerse, Kling, etc. have enabled video generation from free-form textual prompts and demonstrated impressive visual fidelity and alignment with the prompts. However, as illustrated in Figure 1, the T2V-generated videos still exhibit several notable deficiencies that cannot be overlooked. These shortcomings can significantly detract from the user’s viewing experience. Therefore, developing a comprehensive and effective text-to-video quality assessment strategy is essential to ensure an optimal user experience and to guide the optimization of T2V models.
@@ 2.1 Subjective evaluation and dataset for T2V Generation
Chivileva et al. Chivileva et al. (2023) constructed the first T2V generated video evaluation dataset. They obtained 201 textual prompts by combining ChatGPT and manual writing, covering topics such as famous figures, regions, and special festivals. Five T2V models, namely Aphantasia, Text2Video-Zero, T2VSynthesis, Tune-a-Video, and Video Fusion, were selected to generate 1005 videos. Subsequently, 24 subjects were invited to subjectively score the generated videos from the perspectives of naturalness and consistency. Liu et al. proposed FETV Liu et al. (2024b). They designed a set of diverse and fine-grained textual prompts and classified them into three aspects: main content, control attributes, and text complexity. At the same time, a time category was specifically designed for text-to-video, covering spatial and temporal dimensions. On this basis, the authors selected four T2V models, including CogVideo, Text2Video-zero, ModelScopeT2V, and ZeroScope, and conducted subjective evaluations in terms of static quality, temporal quality, overall alignment, and fine-grained alignment. EvalCrafter Liu et al. (2024a) first constructed a comprehensive textual prompt corpus covering daily objects, attributes, and actions. It specified the metadata of common elements in the real world and expanded them using large language models like ChatGPT to obtain complete text prompts. It also enriched and expanded the prompts through those written by real-world users and from text-to-image (T2I) datasets. Subsequently, 2500 videos were generated on 5 T2V models, and subjective evaluations were carried out in terms of video quality, text-video consistency, motion quality, and time consistency. Wu et al. Wu et al. (2024b) constructed the TVGE dataset. They adopted the prompt set proposed in EvalCrafter Liu et al. (2024a) and generated 2543 videos on 5 T2V models. Ten professional scorers were invited to score the videos in terms of video quality and video-text consistency. Huang et al. Huang et al. (2024) believed that the existing evaluation benchmarks could not accurately reflect the highlights and problems of different generation models. Based on this, they proposed Vbench, which further refined the evaluation dimensions of T2V. Sixteen fine-grained evaluation dimensions for generated videos were designed, including subject identity inconsistency, motion smoothness, time flicker, and spatial relationship. One hundred text prompts were designed for each evaluation dimension. In addition, the authors also evaluated T2V models for different content categories, such as people, animals, and landscapes. A total of 8 content categories were considered, and 100 prompts were designed for each category. Based on this, the authors conducted subjective evaluations on 4 T2V models and collected corresponding 16-dimensional scores. Li et al. Zhang et al. (2024c) constructed the LGVQ dataset, which contains 468 text prompts covering combinations of different foreground contents, background contents, and motion states. Six T2V models were used to generate 2808 videos, and 60 subjects were invited to perceptually score the spatial quality, temporal quality, and text-video alignment of each video. Based on the LGVQ dataset, the authors carried out benchmark tests using 14 IQA methods, 16 VQA methods, and 9 methods based on CLIP and visual question answering. Zhang et al. Zhang et al. (2024b) pointed out that the existing subjective evaluation benchmarks for T2V generally have limitations in reliability, reproducibility, and practicability. In view of this, this paper proposes the T2VHE benchmark, which specifies the evaluation dimensions (including video quality, temporal quality, motion quality, text alignment, etc.) used in T2V evaluation, the annotator training process and example descriptions, and a user-friendly GUI interface. The authors further proposed a dynamic scoring scheme. By combining the automatic scoring method to pre-annotate the videos, the amount of manually annotated data was subsequently reduced by 50%. Kou et al. Kou et al. (2024) believed that the current T2V evaluation benchmarks have overly complex divisions of evaluation dimensions and lack a simple and effective evaluation index. Therefore, the authors constructed T2VQA-DB and invited 24 subjects to score 10,000 text-generated videos. The subjects were required to give an overall score considering video quality and text-video consistency.
@@ 2.2 Objective Metrics for T2V Generation
Chivileva et al. Chivileva et al. (2023) proposed the first T2V evaluation index, which assesses the naturalness of the video and the similarity between the text and the video. In the text similarity evaluation, corresponding captions are generated for each video based on the BLIP model. Then, the similarity between the generated captions and the text prompts is calculated using the text encoder of BLIP. For the evaluation of naturalness, a series of handcrafted features are introduced, including texture, sharpness, color, spectrum, entropy, contrast, keypoint detection, and the number and size of speckles. FETV Liu et al. (2024b) aimed at the problem that previous text-to-video evaluation metrics such as FVD, CLIPSim, and IS could not accurately reflect human subjective perception. It improved FVD and CLIPSim and proposed the index FVD-UMT for evaluating video fidelity and the index UMTScore for evaluating text-video alignment. Wu et al. Wu et al. (2024b) proposed T2VScore to automatically evaluate text-to-video in terms of video quality and text-video consistency. T2VScore consists of two components, T2VScore-A and T2VScore-Q, which evaluate text-video consistency and video quality respectively. Specifically, T2VScore-A uses LLMs to generate question-answer pairs for the prompt and then uses MLLMs to ask questions about the video and answer them. The consistency score is obtained by calculating the answer accuracy. The structure of T2VScore-Q adopts a mixture of experts structure, and the basic architecture adopts Fast-VQA. EvalCrafter Liu et al. (2024a) believes that traditional metrics such as FVD and IS cannot well fit human perception, and simple metrics cannot achieve effective evaluation. The article divides video quality, consistency, motion quality, and time consistency into fine-grained sub-dimensions, obtaining a total of 17 evaluation dimensions, and designs a dedicated evaluation algorithm for each dimension (for example, to evaluate the ability of the current T2V model to generate text, EvalCrafter directly uses the PaddleOCR toolkit Ma et al. (2019) to extract text from the generated video. Then, the word error rate, normalized edit distance, and character error rate are calculated and averaged). VBench Huang et al. (2024) also makes a finer division based on the evaluation of video quality and text-video consistency, finally obtaining 16 evaluation dimensions and corresponding 16 evaluation algorithms. Kou et al. Kou et al. (2024) designed the first end-to-end text-to-video evaluation model T2VQA based on the proposed T2VQA-DB dataset. It includes a video fidelity encoding module based on Siwn-3D and a text-video consistency representation module based on BLIP. Finally, the quality score is regressed through a large language model. Lu et al. Lu et al. (2024) improved on T2VQA by introducing a progressive training strategy, enabling different branches of the model to gradually learn relevant knowledge and achieving SOTA performance on T2VQA-DB. Li et al. Zhang et al. (2024c) found that the current metrics could not comprehensively and effectively evaluate text-to-video based on the LGVQ benchmark test. Therefore, the authors proposed the UGVQ model, which systematically extracts spatial, temporal, and text-video alignment features by combining the ViT model, SlowFast network, and CLIP model, and obtains multi-dimensional scores through MLP regression.
@@ 3.1 Division of Subjective Evaluation Dimensions
  • Realness: Whether the video appears to be machine-generated. For videos depicting real-world scenes, this dimension evaluates their conformity with reality; for surreal scenes, it assesses how well the video matches the evaluator’s expectation of animated, sci-fi, or other surreal content.
@@ 3.2 Compilation and Screening of Prompts
Table 1: Detailed information about the models included in this article, including frame rate, duration, resolution, abilities and release date of the version used in the experiment. *Sora has not yet opened its interface, and only 48 published videos participated in our evaluation.
@@ 3.4 Collection of Subjective Scores
We use 95% confidence intervals to detect and remove outliers from the raw scores. For the score of a video generated by model $i$ for prompt $j$ in dimension $k$, the confidence interval is calculated as $[\mu_{ijk}-2\sigma_{ijk},\mu_{ijk}+2\sigma_{ijk}]$. Any scores falling outside this interval are considered outliers and removed.
@@ 4.4 Integrated Scoring prediction Module
LLM has achieved excellent results on IQA and VQA tasks Wu et al. (2024a, 2023b, 2025); Zhang et al. (2024d). Inspired by this, we employed LLM as the scoring prediction head in T2VEval and guided LLM by designing a special prompt "Please rate the quality of the video". The output of LLM is the probability distribution of the five quality levels (bad, poor, fair, good, and excellent). We assigned weights of 1-5 to these five quality levels. Finally, the predicted score was obtained by calculating the softmax of each token and multiplying it by the corresponding weight.
@@ 4.5 Training Strategy
$L~=~L_{plcc}~+~\lambda~\cdot~L_{rank}$ (11)
@@ 5.2 Performance Comparison
Additionally, we fine-tuned and tested T2VEval on the T2VQA-DB dataset, following the same data partitioning method as T2VEval-Bench. The results, averaged across 10 partitions, are also presented in Table 3. These experiments confirm that T2VEval also achieves state-of-the-art performance in this dataset.

### v7

@@ 1 Introduction
Text-to-video (T2V) technology has made significant advancements in the last two years and attracted growing attention from both the research community and the general public. As an important branch of Generative Artificial Intelligence (GenAI), T2V models Blattmann et al. (2023b, a); Ho et al. (2022); Singer et al. (2022); Wu et al. (2021, 2022a); Hong et al. (2022); Villegas et al. (2022); Zhou et al. (2022); Khachatryan et al. (2023); Luo et al. (2023); He et al. (2022); Wang et al. (2023); Yang et al. (2024); Zhang et al. (2024a); Wang et al. (2024b); Brooks et al. (2024) have developed rapidly recently. Contemporary T2V models, such as Sora, Runway Gen-3, PixVerse, Kling, etc. have enabled video generation from free-form textual prompts and demonstrated impressive visual fidelity and alignment with the prompts. However, as illustrated in Figure 1, the T2V-generated videos still exhibit several notable deficiencies that cannot be overlooked. These shortcomings can significantly detract from the user’s viewing experience. Therefore, developing a comprehensive and effective text-to-video quality assessment strategy is essential to ensure an optimal user experience and to guide the optimization of T2V models.
@@ 2.1 Subjective evaluation and dataset for T2V Generation
Chivileva et al. Chivileva et al. (2023) constructed the first T2V generated video evaluation dataset. They obtained 201 textual prompts by combining ChatGPT and manual writing, covering topics such as famous figures, regions, and special festivals. Five T2V models, namely Aphantasia, Text2Video-Zero, T2VSynthesis, Tune-a-Video, and Video Fusion, were selected to generate 1005 videos. Subsequently, 24 subjects were invited to subjectively score the generated videos from the perspectives of naturalness and consistency. Liu et al. proposed FETV Liu et al. (2024b). They designed a set of diverse and fine-grained textual prompts and classified them into three aspects: main content, control attributes, and text complexity. At the same time, a time category was specifically designed for text-to-video, covering spatial and temporal dimensions. On this basis, the authors selected four T2V models, including CogVideo, Text2Video-zero, ModelScopeT2V, and ZeroScope, and conducted subjective evaluations in terms of static quality, temporal quality, overall alignment, and fine-grained alignment. EvalCrafter Liu et al. (2024a) first constructed a comprehensive textual prompt corpus covering daily objects, attributes, and actions. It specified the metadata of common elements in the real world and expanded them using large language models like ChatGPT to obtain complete text prompts. It also enriched and expanded the prompts through those written by real-world users and from text-to-image (T2I) datasets. Subsequently, 2500 videos were generated on 5 T2V models, and subjective evaluations were carried out in terms of video quality, text-video consistency, motion quality, and time consistency. Wu et al. Wu et al. (2024b) constructed the TVGE dataset. They adopted the prompt set proposed in EvalCrafter Liu et al. (2024a) and generated 2543 videos on 5 T2V models. Ten professional scorers were invited to score the videos in terms of video quality and video-text consistency. Huang et al. Huang et al. (2024) believed that the existing evaluation benchmarks could not accurately reflect the highlights and problems of different generation models. Based on this, they proposed Vbench, which further refined the evaluation dimensions of T2V. Sixteen fine-grained evaluation dimensions for generated videos were designed, including subject identity inconsistency, motion smoothness, time flicker, and spatial relationship. One hundred text prompts were designed for each evaluation dimension. In addition, the authors also evaluated T2V models for different content categories, such as people, animals, and landscapes. A total of 8 content categories were considered, and 100 prompts were designed for each category. Based on this, the authors conducted subjective evaluations on 4 T2V models and collected corresponding 16-dimensional scores. Li et al. Zhang et al. (2024d) constructed the LGVQ dataset, which contains 468 text prompts covering combinations of different foreground contents, background contents, and motion states. Six T2V models were used to generate 2808 videos, and 60 subjects were invited to perceptually score the spatial quality, temporal quality, and text-video alignment of each video. Based on the LGVQ dataset, the authors carried out benchmark tests using 14 IQA methods, 16 VQA methods, and 9 methods based on CLIP and visual question answering. Zhang et al. Zhang et al. (2024b) pointed out that the existing subjective evaluation benchmarks for T2V generally have limitations in reliability, reproducibility, and practicability. In view of this, this paper proposes the T2VHE benchmark, which specifies the evaluation dimensions (including video quality, temporal quality, motion quality, text alignment, etc.) used in T2V evaluation, the annotator training process and example descriptions, and a user-friendly GUI interface. The authors further proposed a dynamic scoring scheme. By combining the automatic scoring method to pre-annotate the videos, the amount of manually annotated data was subsequently reduced by 50%. Kou et al. Kou et al. (2024) believed that the current T2V evaluation benchmarks have overly complex divisions of evaluation dimensions and lack a simple and effective evaluation index. Therefore, the authors constructed T2VQA-DB and invited 24 subjects to score 10,000 text-generated videos. The subjects were required to give an overall score considering video quality and text-video consistency. Considering that current AIGVs usually have unique distortions, such as unrealistic objects, unnatural movements, or inconsistent visual elements, it is particularly important to accurately evaluate the quality of AIGVs. Wang et al. Wang et al. (2024a) constructed a large-scale dataset, AIGVQA-DB, which includes 36,576 AIGVs generated by 15 advanced text-to-video models using 1,048 different prompt words. Collect subjective scores from four aspects: static quality, temporal smoothness, dynamic degree, and text-video consistency, and provide pairwise comparison results between videos.
@@ 2.2 Objective Metrics for T2V Generation
Chivileva et al. Chivileva et al. (2023) proposed the first T2V evaluation index, which assesses the naturalness of the video and the similarity between the text and the video. In the text similarity evaluation, corresponding captions are generated for each video based on the BLIP model. Then, the similarity between the generated captions and the text prompts is calculated using the text encoder of BLIP. For the evaluation of naturalness, a series of handcrafted features are introduced, including texture, sharpness, color, spectrum, entropy, contrast, keypoint detection, and the number and size of speckles. FETV Liu et al. (2024b) aimed at the problem that previous text-to-video evaluation metrics such as FVD, CLIPSim, and IS could not accurately reflect human subjective perception. It improved FVD and CLIPSim and proposed the index FVD-UMT for evaluating video fidelity and the index UMTScore for evaluating text-video alignment. Wu et al. Wu et al. (2024b) proposed T2VScore to automatically evaluate text-to-video in terms of video quality and text-video consistency. T2VScore consists of two components, T2VScore-A and T2VScore-Q, which evaluate text-video consistency and video quality respectively. Specifically, T2VScore-A uses LLMs to generate question-answer pairs for the prompt and then uses MLLMs to ask questions about the video and answer them. The consistency score is obtained by calculating the answer accuracy. The structure of T2VScore-Q adopts a mixture of experts structure, and the basic architecture adopts Fast-VQA. EvalCrafter Liu et al. (2024a) believes that traditional metrics such as FVD and IS cannot well fit human perception, and simple metrics cannot achieve effective evaluation. The article divides video quality, consistency, motion quality, and time consistency into fine-grained sub-dimensions, obtaining a total of 17 evaluation dimensions, and designs a dedicated evaluation algorithm for each dimension (for example, to evaluate the ability of the current T2V model to generate text, EvalCrafter directly uses the PaddleOCR toolkit Ma et al. (2019) to extract text from the generated video. Then, the word error rate, normalized edit distance, and character error rate are calculated and averaged). VBench Huang et al. (2024) also makes a finer division based on the evaluation of video quality and text-video consistency, finally obtaining 16 evaluation dimensions and corresponding 16 evaluation algorithms. Kou et al. Kou et al. (2024) designed the first end-to-end text-to-video evaluation model T2VQA based on the proposed T2VQA-DB dataset. It includes a video fidelity encoding module based on Siwn-3D and a text-video consistency representation module based on BLIP. Finally, the quality score is regressed through a large language model. Lu et al. Lu et al. (2024) improved on T2VQA by introducing a progressive training strategy, enabling different branches of the model to gradually learn relevant knowledge and achieving SOTA performance on T2VQA-DB. Li et al. Zhang et al. (2024d) found that the current metrics could not comprehensively and effectively evaluate text-to-video based on the LGVQ benchmark test. Therefore, the authors proposed the UGVQ model, which systematically extracts spatial, temporal, and text-video alignment features by combining the ViT model, SlowFast network, and CLIP model, and obtains multi-dimensional scores through MLP regression. Wang et al. Wang et al. (2024a) proposed AIGV-Assessor, which utilizes spatio-temporal features and the LMM framework to capture the complex quality attributes of AIGVs, thus accurately predicting precise video quality scores and video pair preferences.
In addition, there has been significant progress in the evaluation of Text-to-Image (T2I) generation. To address the remaining issues related to perceptual quality and text-image alignment in generated images, Wang et al. Wang et al. (2025b) proposed a comprehensive dataset and a benchmark, EvalMi-50K, for evaluating large-scale multimodal image generation. Based on EvalMi-50K, they further proposed LMM4LMM, an LMM-based metric method for evaluating large-scale multimodal T2I generation from multiple dimensions, including perception, text-image correspondence, and task-specific accuracy. Wang et al. Wang et al. (2025a) introduce AIGCIQA2023+, a large-scale database for AI-generated image quality assessment with multi-perspective human preference scores and detailed explanations. They propose MINT-IQA, a vision-language model that uses instruction tuning to assess and explain image quality, authenticity, and text-image correspondence, achieving state-of-the-art results on AIGC and traditional IQA benchmarks.
2.3 Video and Image Quality Assessment
Considerable and remarkable progress has been made in the research on traditional video and image quality assessment. Typical works such as Min et al. Min et al. (2017) proposed a blind image quality assessment method based on pseudo-reference images (BPRI). By comparing the structural similarity between a distorted image and its extremely degraded version, BPRI achieves quality evaluation without the need for subjective score training. This method performs well across various databases and is applicable to multiple distortion types and content scenarios. Min et al. Min et al. (2018) proposed the BMPRI method, which generates multiple pseudo-reference images by further degrading the distorted image in various ways and compares their structural similarities. Zhai and Min et al. Zhai Guang Tao (2020) provide a comprehensive survey of perceptual image quality assessment (IQA), reviewing major subjective IQA databases and classifying objective IQA methods, including both traditional and emerging approaches. The paper also compares state-of-the-art IQA algorithms and discusses current challenges and future trends in the field. Min et al. Min et al. (2021) present a comprehensive review of screen content quality assessment, highlighting the unique characteristics and challenges of computer-generated screen content compared to natural scenes. The paper surveys existing methodologies, benchmarks, and evaluation results, and discusses open issues and future research directions in this emerging field. Zhu et al. Zhu et al. (2023) propose a blind image quality assessment model based on contrastive learning and cross-view consistency, integrating color and pseudo-reference information with a Transformer-based feature fusion module. The method achieves strong performance on in-the-wild IQA datasets, demonstrating improved generalization to authentic distortions. Min et al. Min et al. (2024a) present an up-to-date survey of perceptual video quality assessment, summarizing key subjective and objective methods, major databases, and recent advances, including deep learning-based techniques and new application areas. Zhang et al. Zhang et al. (2024c) introduce UGC-New, a large-scale UGC video quality database comprising 405 original and 2430 compressed videos, with extensive subjective ratings. The authors benchmark various VQA models on this dataset and propose an improved model, New-VQA, for better quality assessment of compressed UGC videos. Min et al. Min et al. (2024b) proposed the RichIQA method, which, for the first time, leverages rich subjective information—including mean opinion score (MOS), standard deviation of opinion scores (SOS), and distribution of opinion scores (DOS)—for image quality assessment in the wild. By adopting a three-stage network architecture and multi-label joint training, RichIQA significantly improves the modeling of subjective diversity and generalization across databases. Sun et al. Sun et al. (2025) proposed E-VQA, an efficient video quality assessment model based on deep learning that can strike a balance between high accuracy and low computational cost. Through a systematic analysis of the training components and extensive experiments conducted on user-generated content (UGC) datasets (KVQ, LSVQ), E-VQA demonstrated strong performance and practical efficiency, ranking third in the 2025 NTIRE UGC Video Quality Assessment Challenge.
@@ 3.1 Division of Subjective Evaluation Dimensions
  • Realness: Any content in generative videos that violates human common sense and the physical laws of the real world is not allowed. This indicator mainly quantifies whether there are situations in the video that violate physical laws (such as water not splashing when falling on the ground), physiological laws (a person having six fingers), and any human common sense (such as a sudden change in the number of subjects, inconsistent identity IDs, etc.).
@@ 3.2 Compilation and Screening of Prompts
Table 1: Detailed information about the models included in this article, including frame rate, duration, resolution, abilities and rease date of the version used in the experiment. *Since Sora did not open a public interface during the experiment, this paper only selects the 48 demos initially released by it for evaluation. The test of the public version of Sora will be presented in subsequent work.
@@ 3.4 Collection of Subjective Scores
We use 95% confidence intervals to detect and remove outliers from the raw scores. For the score of a video generated by model $i$ for prompt $j$ in dimension $k$, the confidence interval is calculated as $\mu_{ijk}-2\sigma_{ijk},\mu_{ijk}+2\sigma_{ijk}]$. Any scores falling outside this interval are considered outliers and removed.
@@ 4.4 Integrated Scoring prediction Module
LLM has achieved excellent results on IQA and VQA tasks Wu et al. (2024a, 2023b, 2025); Zhang et al. (2024e). Inspired by this, we employed LLM as the scoring prediction head in T2VEval and guided LLM by designing a special prompt "Please rate the quality of the video". The output of LLM is the probability distribution of the five quality levels (bad, poor, fair, good, and excellent). We assigned weights of 1-5 to these five quality levels. Finally, the predicted score was obtained by calculating the softmax of each token and multiplying it by the corresponding weight.
@@ 4.5 Training Strategy
$L\nobreakspace=\nobreakspace L_{plcc}\nobreakspace+\nobreakspace\lambda\nobreakspace\cdot\nobreakspace L_{rank}$ (11)
@@ 5.2 Performance Comparison
In addition, to verify the effectiveness and advancement of T2VEval, we fine-tuned and tested T2VEval on the T2VQA-DB, LGVQ, and Q-Eval datasets. Table 3 shows the experimental results on T2VQA-DB, where T2VEval also achieved state-of-the-art performance on this dataset. Table 4 shows the performance of T2VEval on LGVQ. It can be seen that T2VEval achieved the best performance in both the temporal quality and consistency dimensions, but was slightly insufficient in spatial quality. This may be because our backbone networks preferentially adopt 3D spatio-temporal modeling networks, which also takes into account the biggest difference between videos and images: videos are dynamic and change over time. Table 5 shows the performance of our method on Q-Eval, and our method also achieved state-of-the-art performance. It should be noted that since the complete test set of Q-Eval cannot be obtained, we can only perform dataset partitioning and testing on the publicly available training set.
Table 4: Performance comparison on LGVQ. Red: the best
Aspects Methods SROCC PLCC KROCC
Spatial MUSIQ 0.669 0.682 0.491
StairIQA 0.701 0.737 0.521
CLIP-IQA 0.684 0.709 0.502
LIQE 0.721 0.752 0.538
UGVQ 0.759 0.795 0.567
AIGV-Assessor 0.803 0.819 0.617
T2VEval 0.753 0.789 0.560
Temporal VSFA 0.841 0.857 0.643
SimpleVQA 0.857 0.867 0.659
FastVQA 0.849 0.843 0.647
DOVER 0.867 0.878 0.672
UGVQ 0.893 0.907 0.703
AIGV-Assessor 0.900 0.902 0.717
T2VEval 0.906 0.914 0.720
Alignment CLIPScore 0.446 0.453 0.301
BLIPScore 0.455 0.464 0.319
ImageReward 0.498 0.499 0.344
PickScore 0.501 0.515 0.353
HPSv2 0.504 0.511 0.357
UGVQ 0.551 0.555 0.394
AIGV-Assessor 0.577 0.578 0.411
T2VEval 0.674 0.694 0.496
Table 5: Performance comparison on Q-Eval. Red: the best
Aspects Methods PLCC SROCC
Visual Quality SimpleVQA 0.346 0.385
DOVER 0.506 0.506
Q-Align 0.416 0.460
T2VQA 0.516 0.516
T2VEval 0.535 0.517
Alignment CLIPScore 0.443 0.431
BLIP2Score 0.488 0.483
ImageReward 0.485 0.472
T2VEval 0.685 0.688